Solve 11sin^2 y=13-sin^2 y for y, where 0<_y<360. How many angles satisffy this equation? Round anzwer to nearest tenth. what second quadrant satisfies this equation
what fourth quadrant angle satisfy this equation?

Answer by venugopalramana(3286)

Solve 11sin^2 y=13-sin^2 y for y, where 0<_y<360. How many angles satisffy this equation? Round anzwer to nearest tenth. what second quadrant satisfies this equation
Solve 11sin^2 y=13-sin^2 y
12 SIN^2 Y=13
SIN^2Y=13/12.....THIS IS NOT POSSIBLE...SINCE SIN(Y) CAN NEVER BE MORE THAN 1....SO NO SOLUTION.CHECK YOUR PROBLEM AGAIN